Need to know if I would add subtract multiply or divide these numbers to get the right answer it is one number
5 y equals 11
10 y equals 21
15 y equals 31
20 y equals 41
Also same for the next question
14 y equals 2
10 y equals -2
6 y equals -6
2 y equals -10
This is for my daugther If you could please send the answer and explain to me how u got that answer so it will help me moving forward with her homework when she has questions Thank you in advance

Answer by richwmiller(17219)

It is not possible that it is the same number for each of these
5y equals 11 5y=11 y=11/5 y=2.2
10y equals 21 10y=21 y=21/10 y=2.1
The numbers are close but not the same.